OpenStudy (anonymous):

solve for n. 1/n-4 - 2/n = 3/4-n n = 2 n = -2 n = -4 n = 4

OpenStudy (anonymous):

Is that 1/(n-4) - 2/n = 3/(4-n) ?

OpenStudy (anonymous):

yes sorry

OpenStudy (anonymous):

Alrighty that's fine.

OpenStudy (anonymous):

First you could take a negative out of n-4 to make it -(4-n) Then you could multiply everything by n * (4-n) to get rid of all denominators.

OpenStudy (anonymous):

So you would get -n - 2(4-n) = 3n And that is much easier to solve.

OpenStudy (anonymous):

distributing you get -n -8 + 2n = 3n then we collect n's and constants 2n = -8 n = -4